OpenStudy (anonymous):

I NEED HELP Starting from 1.5 miles away, a car drives toward a speed checkpoint and then passes it. The car travels at a constant rate of 53 miles per hour. The distance of the car from the checkpoint is given by d = |1.5 – 53t|. At what times is the car 0.1 miles from the checkpoint? Calculate your answer in seconds.

hero (hero):

What the...See, this is why you should have stayed in your question bro. It was already answered minutes after you left the first time.

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

awww im sorry

hero (hero):

Look over the responses carefully. The steps are there

hero (hero):

d = |1.5 – 53t| 0.1 = -(1.5 - 53t) 0.1 = 1.5 - 53t 0.1 =53t - 1.5 1.5 + 0.1 = 53t 1.6/53 = t t = 108.67 secs 53t = 1.5 - 0.1 53t = 1.4 t = 1.4/53 t = 95.09 secs

hero (hero):

What about it do you need explained?